My dad enjoyed jogging at night. He would have his little MP3 player with him listening to either R&B, jazz, or news while he did this nightly ritual. In a hypothetical situation, Dad had exactly 14 hours and 5 minutes of audio content on his device. The linear equation 6x+4y+125=845 can help determine how many jazz songs, x, and how many R&B songs, y, Dad had on his mp3 player. Based on this equation, how many R&B songs and jazz songs combined were on my dad's device if the number of songs for each genre was the same?
Answer by solve_for_x(190)

If the number of R&B songs and the number of jazz songs are the same, then x = y.

Substituting y in place of x in the equation gives:

6x + 4x + 125 = 845

10x + 125 = 845

10x = 845 - 125

10x = 720

x = 72

Solution: Dad has 72 R&B songs and 72 jazz songs on his mp3 player. Combined, he has 144 R&B and jazz songs.